# An approximate Kappa generator for particle simulations

### v1

@@ IV Random Number generator
$\ln_{q}(x)\equiv\dfrac{x^{1-q}-1}{1-q}.$ (23)
One can solve an inverse problem of Eq. (22), by way of a quadratic equation for xx.(Umeda, 2024) Then we obtain the particle velocity $v$ in the following way:
$v$ $=\theta\sqrt{\frac{-(a+cL)+\sqrt{(a+cL)^{2}-4bL}}{2b}}$
$=\theta\sqrt{\frac{-2L}{(a+cL)+\sqrt{(a+cL)^{2}-4bL}}}$ (24)
@@
$b\leftarrow\left(\kappa^{*}\frac{3}{2}B(3/2,\kappa^{*})\right)^{1/\kappa^{*}}\frac{\kappa^{*}}{\kappa}\,c$
@@ V Numerical tests
To validate the proposed method in Table 1, we have carried out three numerical tests. If necessary, we compare the proposed method to two known methods: (1) the standard Kappa generator, based on the $t$-distribution,(Abdul & Mace, 2015; Zenitani & Nakano, 2022) and (2) the rejection method, based on the Pareto distribution.(Zenitani, 2025) We call them the standard method and the Pareto method, respectively. Inside the standard method, we employ the Marsaglia & Tsang (2000) method to generate a gamma variate. The Pareto method has two options in its index $n$, but we choose a simple option $n=\kappa/2$, because it runs substantially faster than the other option.
@@
Figure 6 shows the relative entropy $D$ as a function of the total particle number for $\kappa=3.0$ (red), $4.1$ (orange), and $7.5$ (blue). Monte Carlo results are calculated by two methods. The dashed lines indicate the results by the standard $t$-generator method,(Abdul & Mace, 2015; Zenitani & Nakano, 2022) while the solid lines indicate those by the proposed method. To clarify, the standard method essentially draws random numbers from the exact Kappa distribution, while the proposed method draws random numbers from the approximated distribution.
@@
Third, finally, we have evaluated the performance of the three Kappa generators: (1) the standard method, (2) the Pareto method, and (3) the proposed method. We wrote our CPU codes in C, compiled them with the clang compiler (v18.1) with the -lgsl -lm options, and ran them on AMD Ryzen 5955 processor on Ubuntu Linux 24.04LTS. Similarly, we wrote our GPU codes in CUDA, compiled them with the nvcc compiler (v12.0), and ran them with an NVIDIA RTX A6000 GPU. We carried out a series of runs to generate $10^{8}$ particles, as a function of $\kappa$, ranging from 1.6 to 15. For each case, we ran the code four times, and then averaged elapsed times of the last three runs.
@@
Figure 7 shows the results. On the CPU (Figure 7(a)), we find that the proposed method (in gray) runs faster than the standard method (in blue). Furthermore, the Pareto method is the fastest, possibly due to its simplicity. The Pareto method shows weak $\kappa$-dependent variation, while the other methods show only little variation. On the GPU (Figure 7(b)), the proposed method runs substantially faster than the other methods. It essentially shows no variation, regardless of $\kappa$. In contrast, the other methods show significant variation. We attribute these trends to the nature of the SIMT algorithms, as discussed in the next section.
@@ VI Discussion and Summary
In the first and second tests, we find that the proposed method provides a remarkably good approximation. The results are visually indistinguishable from the exact solutions (Fig. 5), as the difference around the most probable speed $v\approx\theta$ is kept small (Fig. 3(b)). In the low-$\kappa$ regime ($\kappa<4$), the proposed method gives a near-perfect approximation, as highlighted in the low $\mathcal{D}$ values (Fig. 4(a)). The Monte Carlo results with $\kappa=3.0$ are indistinguishable from the Kappa distribution in Fig. 6. Extrapolating the results toward the asymptotic value, we will see a difference when $N_{p}\gtrsim 10^{9.5}$. In the high-$\kappa$ regime ($\kappa\gtrsim 5$), the difference between the two distributions emerges in the high-energy part (Fig. 3(b)). Since phase space density in the high-energy part is already very low for high-$\kappa$ cases, the overall properties of the distribution are unlikely to be changed. Figures 4(a) and 4(b) tell us that the discrepancy is most pronounced in the transition case of $\kappa\approx 4.1$. In Fig. 6, the relative entropies deviate when $N_{p}\gtrsim 10^{7.5}$ ($\kappa\gtrsim 5$) and $N_{p}\gtrsim 10^{7}$ ($\kappa\approx 4.1$). Below these particle numbers, the difference between the approximate and exact Kappa distributions is completely hidden by the particle noise. Since today’s PIC simulation typically uses $10^{2-3}$ particles in a grid cell, the proposed method would be accurate enough in many cases. The difference between the two distributions certainly emerges in $D$ beyond these numbers, however, even in the large-number limit, the PSD profile still looks similar to that of the Kappa distribution (Fig. 3(a)), at least on the logarithmic scale. Probably the biggest concern is the total energy density (Fig. 4(b)). The normalized error is within $10^{-3}$ in both low- and high-$\kappa$ regimes, but it increases to $\mathcal{O}(10^{-2.5})$ in the transition case. If the user wants to explore the energy balance at the accuracy of $\mathcal{O}(10^{-2.5})$, the proposed method is not ideal.
@@
For simplicity, we estimate the computational costs by using the total number of the uniform or normal variates per particle. All the other factors are ignored. Table 1 tells us that the proposed method requires three uniform variates. The standard method calls three normal variates, and two (one uniform and one normal) variates inside the gamma RNG, which employs the Marsaglia & Tsang (2000)’s rejection method (hereafter the MT method). The acceptance rate of the MT method is given by (Ref. Yotsuji, 2010, eq. 3.308),
@@
The situation is different on GPUs. Inside GPUs, typically 32 threads are controlled by the same instruction. In such a case, one can proceed to the next computational step, only after all 32 threads complete the loop. According to Ridley & Forget (2021), the typical count $\mu$ of the rejection loop is approximated by the mode of the exponentiated geometric distribution,(Chacrabaty & Gupta, 2015)
@@
In summary, we have proposed a simple inverse transform method for generating the Kappa distribution in particle simulations. The method is fully described in Table 1. It generates the approximate distribution, which is particularly close to the Kappa distribution in the low-$\kappa$ regime. The relation between the particle number, accuracy, and $\kappa$ parameters is discussed. Computationally, the new method outperforms the conventional methods in many cases on the CPU and on the GPU. By definition, the proposed method is GPU-friendly, as it is free from the side effects in the SIMT model. We hope that the proposed method will be useful on future GPUs with more SIMT threads.
@@ Acknowledgements
TU was supported by Grant-in-Aid for Scientific Research (B) 24K00603 from the Japan Society for the Promotion of Science (JSPS).

### v2

@@ IV Random Number generator
$\ln_{q}(x)\equiv\dfrac{x^{1-q}-1}{1-q},$ (23)
one can solve an inverse problem of Eq. (22), by way of a quadratic equation for $x$ (Umeda, 2024). Then we obtain the particle velocity $v$ in the following way:
$v$ $=\theta\sqrt{\frac{-(a+cL)+\sqrt{(a+cL)^{2}-4bL}}{2b}}=\theta\sqrt{\frac{-2L}{(a+cL)+\sqrt{(a+cL)^{2}-4bL}}}$ (24)
@@
$b\leftarrow\left\{\kappa^{*}\frac{3}{2}B(3/2,\kappa^{*})\right\}^{1/\kappa^{*}}\frac{\kappa^{*}}{\kappa}\,c$
@@ V Numerical tests
To validate the proposed method in Table 1, we have carried out three numerical tests. If necessary, we compare the proposed method to two known methods: (1) the standard Kappa generator, based on the $t$-distribution (Abdul & Mace, 2015; Zenitani & Nakano, 2022), and (2) the rejection method, based on the Pareto distribution (Zenitani, 2025). We call them the standard method and the Pareto method, respectively. Inside the standard method, we employ the Marsaglia & Tsang (2000) method to generate a gamma variate. The Pareto method has two options in its index $n$, but we choose a simple option $n=\kappa/2$, because it runs substantially faster than the other option.
@@
Figure 6 shows the relative entropy $D$ as a function of the total particle number for $\kappa=3.0$ (red), $4.1$ (orange), and $7.5$ (blue). Monte Carlo results are calculated by two methods. The dashed lines indicate the results by the standard $t$-generator method (Abdul & Mace, 2015; Zenitani & Nakano, 2022), while the solid lines indicate those by the proposed method. To clarify, the standard method essentially draws random numbers from the exact Kappa distribution, while the proposed method draws random numbers from the approximated distribution.
@@
Third, finally, we have evaluated the performance of the three Kappa generators: (1) the standard method, (2) the Pareto method, and (3) the proposed method. We wrote our CPU codes in C, compiled them with the clang compiler (v18.1) with the -lgsl -lm options, and ran them on AMD Ryzen 5955 processor on Ubuntu Linux 24.04LTS. Similarly, we wrote our GPU codes in CUDA, compiled them with the nvcc compiler (v12.0), and ran them with an NVIDIA RTX A6000 GPU. We carried out a series of runs to generate $10^{8}$ particles, as a function of $\kappa$, ranging from 1.6 to 15. To completely rule out startup issues, for each case, we ran the code four times, and then averaged elapsed times of the last three runs.
@@
Figure 7 shows the results. On the CPU (Figure 7(a)), we find that the proposed method (in gray) runs faster than the standard method (in blue). Furthermore, the Pareto method is the fastest, possibly due to its simplicity. The Pareto method shows a weak $\kappa$-dependent variation, while the other methods show only little variation. On the GPU (Figure 7(b)), the proposed method runs substantially faster than the other methods. It essentially shows no variation, regardless of $\kappa$. In contrast, the other methods show significant variation. We attribute these trends to the nature of the SIMT algorithms, as discussed in the next section.
@@ VI Discussion and Summary
In the first and second tests, we find that the proposed method provides a remarkably good approximation. The results are visually indistinguishable from the exact solutions (Fig. 5), as the difference around the most probable speed $v\approx\theta$ is kept small (Fig. 3(b)). In the low-$\kappa$ regime ($\kappa<4$), the proposed method gives a near-perfect approximation, as highlighted in the low $\mathcal{D}$ values (Fig. 4(a)). The Monte Carlo results with $\kappa=3.0$ are indistinguishable from the Kappa distribution in Fig. 6. Extrapolating the results toward the asymptotic value, we will see a difference when $N_{p}\gtrsim 10^{9.5}$. In the high-$\kappa$ regime ($\kappa\gtrsim 5$), the difference between the two distributions emerges in the high-energy part (Fig. 3(b)). Since phase space density in the high-energy part is already very low for high-$\kappa$ cases, the overall properties of the distribution are unlikely to be changed. Figures 4(a) and 4(b) tell us that the discrepancy is most pronounced in the transition case of $\kappa\approx 4.1$. In Fig. 6, the relative entropies deviate when $N_{p}\gtrsim 10^{7.5}$ ($\kappa\gtrsim 5$) and $N_{p}\gtrsim 10^{7}$ ($\kappa\approx 4.1$). Below these particle numbers, the difference between the approximate and exact Kappa distributions is completely hidden by the particle noise. Since today’s PIC simulation typically uses $10^{2-3}$ particles in a grid cell, the proposed method would be accurate enough in many cases. The difference between the two distributions certainly emerges in $D$ beyond these numbers; however, even in the large-number limit, the PSD profile still looks similar to that of the Kappa distribution (Fig. 3(a)), at least on the logarithmic scale. Probably the biggest concern is the total energy density (Fig. 4(b)). The normalized error is within $10^{-3}$ in both low- and high-$\kappa$ regimes, but it increases to $\mathcal{O}(10^{-2.5})$ in the transition case. If the user wants to explore the energy balance at the accuracy of $\mathcal{O}(10^{-2.5})$, the proposed method is not ideal.
@@
For simplicity, we estimate the computational costs by using the total number of the uniform or normal variates per particle. All the other factors are ignored. Table 1 tells us that the proposed method requires three uniform variates. The standard method calls three normal variates, and two (one uniform and one normal) variates inside the gamma RNG, which employs the Marsaglia & Tsang (2000)’s rejection method (hereafter the MT method). The acceptance rate of the MT method is given by (Yotsuji (2010), eq. 3.308),
@@
The situation is different on GPUs. Inside GPUs, typically 32 threads are controlled by the same instruction. In such a case, one can proceed to the next computational step, only after all 32 threads complete the loop. According to Ridley & Forget (2021), the typical count $\mu$ of the rejection loop is approximated by the mode of the exponentiated geometric distribution (Chacrabaty & Gupta, 2015),
@@
In summary, we have proposed a simple inverse transform method for generating the Kappa distribution in particle simulations. The method is fully described in Table 1. It generates an approximate distribution, which is particularly close to the Kappa distribution in the low-$\kappa$ regime. The relations between the particle number, accuracy, and $\kappa$ parameters are discussed. Computationally, the new method outperforms the conventional methods in many cases on the CPU and on the GPU. By definition, the proposed method is GPU-friendly, as it is free from the side effects in the SIMT model. We hope that the proposed method will be useful on future GPUs with more SIMT threads.
@@ Acknowledgements
TU was supported by Grant-in-Aid for Scientific Research (B) 24K00603 from the Japan Society for the Promotion of Science (JSPS). The Jupyter notebook and C/CUDA codes are archived in Zenodo (Zenitani & Umeda, 2026).
Appendix A Taylor expansions
Here we formally derive Eqs. (11), (12), (14), and (15). We first evaluate the exact CDF of Eq. (3). For convenience, we use a new variable $y\equiv x/(x+\kappa)$. Near $x=0$, we find
$y$ $=\frac{x}{\kappa}-\left(\frac{x}{\kappa}\right)^{2}+\mathcal{O}(x^{3})$ (32)
$y^{3/2}$ $=\left(\frac{x}{\kappa}\right)^{3/2}\left(1-\frac{x}{\kappa}\right)^{3/2}=\left(\frac{x}{\kappa}\right)^{3/2}\left(1-\frac{3}{2\kappa}x+\mathcal{O}(x^{2})\right)$ (33)
Considering the definition of the incomplete beta function, we expand Eq. (3) as follows:
$F(x)$ $=I_{y}\left(\frac{3}{2},\kappa-\frac{1}{2}\right)=\frac{1}{B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\int_{0}^{y}t^{1/2}(1-t)^{\kappa-3/2}dt$
$=\frac{1}{B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\int_{0}^{y}t^{1/2}\bigg(1-(\kappa-3/2)t+\mathcal{O}(t^{2})\bigg)dt$
$=\frac{y^{3/2}}{\frac{3}{2}B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\bigg(1+\frac{3({3/2-\kappa})}{5}y+\mathcal{O}(y^{2})\bigg)$ (34)
We can also verify this expansion, with help from the hypergeometric function $_{2}F_{1}(a,b;c;x)$ and formulas in Olver et al. (2025, 8.1.7, 15.2.1),
$F(x)$ $=\frac{y^{3/2}}{\frac{3}{2}B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}{}_{2}F_{1}\left(\frac{3}{2},\frac{3}{2}-\kappa;\frac{5}{2};y\right)=\frac{y^{3/2}}{\frac{3}{2}B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\left(1+\frac{3(3/2-\kappa)}{5}y+\mathcal{O}(y^{2})\right)$ (35)
Substituting Eqs. (32) and (33) into Eq. (34), and since $y^{2}$ is an order of $\mathcal{O}(x^{2})$ , we find
$F(x)$ $=\frac{2}{3B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\left(\frac{x}{\kappa}\right)^{3/2}\left(1+\frac{3(\frac{3}{2}-\kappa)}{5}\frac{x}{\kappa}+\mathcal{O}(x^{2})\right)\left(1-\frac{3}{2\kappa}x+\mathcal{O}(x^{2})\right)$
$=\frac{2}{3B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\left(\frac{x}{\kappa}\right)^{3/2}\left(1-\frac{3(\kappa+1)}{5\kappa}x+\mathcal{O}(x^{2})\right)$ (36)
in agreement with Eq. (11).
Next we examine the $x\rightarrow\infty$ limit. In this case, we expand a new variable $z$ with respect to $(1/x)$.
$z$ $\equiv 1-y={\kappa}{x}^{-1}-{\kappa^{2}}{x}^{-2}+\mathcal{O}(x^{-3})$ (37)
$z^{\kappa-\frac{1}{2}}$ $=\left(\frac{\kappa}{x}\right)^{\kappa-\frac{1}{2}}\left(1-\frac{\kappa}{x}\right)^{\kappa-\frac{1}{2}}=\left(\frac{\kappa}{x}\right)^{\kappa-\frac{1}{2}}\left(1-\kappa\left(\kappa-\frac{1}{2}\right){x^{-1}}+\mathcal{O}(x^{-2})\right)$ (38)
Using the relation $I_{x}(a,b)=1-I_{1-x}(b,a)$, we find
$F(x)$ $=1-I_{z}\left(\kappa-\frac{1}{2},\frac{3}{2}\right)=1-\frac{1}{B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\int_{0}^{z}t^{\kappa-{3}/{2}}(1-t)^{1/2}dt$
$=1-\frac{1}{B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\int_{0}^{z}t^{\kappa-{3}/{2}}\left(1-\frac{1}{2}t+\mathcal{O}(t^{2})\right)dt$
$=1-\frac{z^{\kappa-\frac{1}{2}}}{\left(\kappa-\frac{1}{2}\right)B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\left(1-\frac{\kappa-\frac{1}{2}}{2\kappa+1}z+\mathcal{O}(z^{2})\right)$ (39)
We can similarly check the results using the formulas (Olver et al., 2025, 8.1.7, 15.2.1),
$F(x)$ $=1-\frac{z^{\kappa-\frac{1}{2}}}{\left(\kappa-\frac{1}{2}\right)B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}{}_{2}F_{1}\left(\kappa-\frac{1}{2},-\frac{1}{2};\kappa+\frac{1}{2};z\right)$
$=1-\frac{z^{\kappa-\frac{1}{2}}}{\left(\kappa-\frac{1}{2}\right)B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\left(1-\frac{\kappa-\frac{1}{2}}{2\kappa+1}z+\mathcal{O}(z^{2})\right)$ (40)
Substituting Eqs. (37) and (38) into Eq. (39), and since $z^{2}$ is an order of $\mathcal{O}(x^{-2})$ , we find
$F(x)$ $=1-\frac{\kappa^{\kappa-1/2}}{\left(\kappa-\frac{1}{2}\right)B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\left(1-\frac{\kappa-\frac{1}{2}}{2\kappa+1}\kappa x^{-1}+\mathcal{O}(x^{-2})\right)\left(1-\left(\kappa-\frac{1}{2}\right){\kappa x^{-1}}+\mathcal{O}(x^{-2})\right)x^{-(\kappa-1/2)}$
$=1-\frac{\kappa^{\kappa-1/2}}{\left(\kappa-\frac{1}{2}\right)B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}x^{-(\kappa-1/2)}\left(1-\frac{\left(\kappa-\frac{1}{2}\right)(\kappa+1)\kappa}{\kappa+\frac{1}{2}}{x^{-1}}+\mathcal{O}(x^{-2})\right)$ (41)
This is in agreement with agreement with Eq. (14).
Next we evaluate the approximate CDF of Eq. (7). We rewrite it by using the $q$-exponential function,
$G(x)=\left\{1-\bigg(1+\frac{1}{\kappa^{*}}\frac{ax+bx^{2}}{1+cx}\bigg)^{-\kappa^{*}}\right\}^{3/2}.$ (42)
We expand the fraction part,
$\frac{ax+bx^{2}}{1+cx}=ax+(b-ac)x^{2}+\mathcal{O}(x^{3})$ (43)
and then we expand the $q$-exponential function
$\bigg(1+\frac{1}{\kappa^{*}}\frac{ax+bx^{2}}{1+cx}\bigg)^{-\kappa^{*}}$ $=\bigg(1+\frac{1}{\kappa^{*}}\left\{ax+(b-ac)x^{2}+\mathcal{O}(x^{3})\right\}\bigg)^{-\kappa^{*}}$
$=1-\bigg(ax+(b-ac)x^{2}+\mathcal{O}(x^{3})\bigg)+\frac{(\kappa^{*}+1)}{2\kappa^{*}}\bigg(ax+(b-ac)x^{2}+\mathcal{O}(x^{3})\bigg)^{2}$
$=1-ax+\bigg(\frac{(\kappa^{*}+1)a^{2}}{2\kappa^{*}}-(b-ac)\bigg)x^{2}+\mathcal{O}(x^{3})$ (44)
Substituting Eq. (44) into Eq. (42), we obtain Eq. (12)
$G(x)$ $=\left\{ax+\bigg((b-ac)-\frac{(\kappa^{*}+1)a^{2}}{2\kappa^{*}}\bigg)x^{2}+\mathcal{O}(x^{3})\right\}^{3/2}$
$=a^{3/2}x^{3/2}\left\{1+\bigg(\frac{b-ac}{a}-\frac{(\kappa^{*}+1)a}{2\kappa^{*}}\bigg)x+\mathcal{O}(x^{2})\right\}^{3/2}$
$=a^{3/2}x^{3/2}\left\{1+\frac{3}{2}\left(\frac{b-ac}{a}-\frac{(\kappa^{*}+1)a}{2\kappa^{*}}\right)x+\mathcal{O}(x^{2})\right\}$ (45)
For $x\rightarrow\infty$, we expand the fraction part with respect to $(1/x)$:
$\frac{ax+bx^{2}}{1+cx}=x\left(\frac{b}{c}+\frac{ac-b}{c^{2}}x^{-1}+\mathcal{O}(x^{-2})\right)$ (46)
Using this, the $q$-exponential function yields
$\bigg(1+\frac{1}{\kappa^{*}}\frac{ax+bx^{2}}{1+cx}\bigg)^{-\kappa^{*}}$ $=\bigg(1+\frac{x}{\kappa^{*}}\left\{\frac{b}{c}+\frac{ac-b}{c^{2}}x^{-1}+\mathcal{O}(x^{-2})\right\}\bigg)^{-\kappa^{*}}$
$=\bigg(\frac{c\kappa^{*}}{bx}\bigg)^{\kappa^{*}}\bigg(1+\frac{c}{b}\left\{\kappa^{*}+\frac{ac-b}{c^{2}}\right\}x^{-1}+\mathcal{O}(x^{-2})\bigg)^{-\kappa^{*}}$
$=\bigg(\frac{c\kappa^{*}}{bx}\bigg)^{\kappa^{*}}\bigg(1-\frac{c\kappa^{*}}{b}\left\{\kappa^{*}+\frac{ac-b}{c^{2}}\right\}x^{-1}+\mathcal{O}(x^{-2})\bigg)$ (47)
Substituting Eq. (47) into Eq. (42), we arrive at Eq. (15)
$G(x)$ $=\left\{1-\bigg(\frac{c\kappa^{*}}{b}\bigg)^{\kappa^{*}}x^{-\kappa^{*}}\bigg(1+\frac{c\kappa^{*}}{b}\left\{\kappa^{*}+\frac{ac-b}{c^{2}}\right\}x^{-1}+\mathcal{O}(x^{-2})\bigg)\right\}^{3/2}$
$=1-\frac{3}{2}\left(\frac{c\kappa^{*}}{b}\right)^{\kappa^{*}}x^{-\kappa^{*}}+\mathcal{O}(x^{-(\kappa^{*}+1)})$ (48)